given the functions: f(x) = x2 - 3x +5 and g (x) = 5x - 1; find f (2)
Answer by nyc_function(2741)

Replace x with 2 for f(x) and simplify.
f(2) = (2)^2 - 3(2) + 5
f(2) = 4 - 6 + 5
f(2) = -2 + 5
f(2) = 3